# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v5

@@ IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE V: Transfer-based attack against SOTA detectors on Synthetic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [wang2020cnn].
@@
Attacks against Deepfake detectors are similar to our attacks. However, Deepfakes mainly manipulate or synthesize faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [li2021exploring], removing facial forgery traces [huang2020fakepolisher, wu2024traceevader] or simply updating perturbations in the frequency domain to improve attack imperceptibility [jia2022exploring] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [huang2020fakepolisher], TraceEvader [wu2024traceevader] and FaceAttacker [jia2022exploring], which are the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table III, FPBA consistently achieves the highest attack success rates across all detectors and datasets, outperforming prior Deepfake-based attacks by margins ranging from 11.5% to 87.5%. FakePolisher performs better than TraceEvader and FaceAttack (albeit still less effective than FPBA), but the former largely reduces the image quality, which is very visible and raises suspicion, as shown in Table XIV and Fig. 8. This is likely because FakePolisher is primarily trained on the face manifold, limiting its ability to reconstruct fine-grained, identity-preserving details in AIGI images.
@@ IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [dong2018boosting] and SVRE [svre], i.e., utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table IV. First, compared with setting (1), the heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial examples only generated by CNNSpot with FPBA get a success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
@@ IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [GramNet] extracts the global texture representation; UnivFD [ojha2023towards] captures semantic feature using pre-trained CLIP [CLIP]; LNP [liu2022detecting] extracts the noise pattern of images; LGrad [tan2023learning] extract gradient information through a pre-trained model; DNF [zhang2023diffusion] leverages diffusion noise features, extracting them through an inverse diffusion process [adm]; FreqNet [tan2024frequency] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [doloriel2024frequency] introduces a frequency-domain mask strategy for data augmentation during training. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table V. Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability between each other. Nevertheless, FPBA still achieves the best adversarial transferability.
@@
Very recently, reconstruction error-based methods, such as DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade] have emerged as competitive approaches for detecting diffusion-generated images. Therefore, we adopt them as victim models in our evaluation. As both methods rely on diffusion models, their performance degrades considerably on GAN-generated data. Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX. We follow [ricker2024aeroblade] to report detection performance using average precision (AP). FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
@@ IV-C2 Attack Cross-Generator Image Detection
One important real-world detection problem is cross-generator image detection, i.e., identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 [sd] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [Midjourney], SD V1.4 [sd], SD V1.5 [sd], ADM [adm], Wukong [wukong] and BigGAN [biggan]. Because Swin-ViT achieves the SOTA results on different subsets [zhu2024genimage], we use it as the surrogate model.
@@ IV-D Adversarial Transferability Analysis
Next, considering that the AIGI detection task often suffers from imbalanced recognition accuracy between real and fake samples (as shown in Table VI), we further investigate the attack success rates of FPBA on real images and fake images separately, as shown in Table X. Low-level feature detectors, including FreqMask, FreqNet, GramNet, LGrad, mainly rely on frequency, texture, or spectral cues. We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples. We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [yanorthogonal]. As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack. In contrast, high-level detectors like UnivFD, which rely on global semantic features, exhibit overall lower transferability. This suggests that models leveraging generalizable semantic representations are inherently more robust. Interestingly, we observe that in UnivFD, fake images are more resistant to adversarial attacks than real ones. We will leave the phenomenon to explore in future work.
@@ IV-E Evaluation against Defense Models
Next, we investigate the attack performance against defense models. As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors. In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [pgd] and TRADES [trades]. However, we found that the training process of AT fails to converge. Both AT methods cause a substantial drop in clean accuracy, often to around 50%, which is equivalent to random guessing. Robust accuracy against PGD attacks also remains low, in some cases collapsing to nearly 0%. This phenomenon can be attributed to the intrinsic data imbalance in AIGI detection: models tend to overfit to low-level synthetic artifacts early in training. When AT is applied, these brittle features collapse quickly, and adversarial examples easily cross the decision boundary for both real and fake classes. As a result, the model fails to learn discriminative and robust representations, producing near-random predictions on clean data and negligible robustness gains.
@@
Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies. Considering that AIGI detection typically [wang2020cnn] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [wang2020cnn, zhu2024genimage] to utilize Gaussian blurring and JPEG compression as a defense. In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is also crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 7. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under real-world scenarios.
Next, we follow the suggestions from [wang2020cnn, zhu2024genimage] to simultaneously utilize Gaussian blurring and JPEG compression as a defense. As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@ IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have a higher success rate [dong2018boosting]. It indicates the possibility of gradient masking [athalye2018obfuscated]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 9 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
@@ IV-G2 Evaluation on Deepfake Datasets
As we mentioned before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [shamshad2023evading], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [ffhq]. We report the results in Table XIII. Although FPBA is not specifically designed for Deepfake detection, it still achieves the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
@@ Acknowledgments
[Uncaptioned image] Yunfeng Diao is a Lecturer in the School of Computer Science and Information Engineering, Hefei University of Technology, China. He received his PhD from Southwest Jiaotong University, China. His current research interests include computer vision and the security of machine learning. He has published over 30 papers in leading venues such as CVPR, ICLR, ICCV, ICML, IEEE TMM and IEEE TCSVT. He has been an editorial board member of IJAACS, the chief organizer of several workshops at IJCAI, also regularly reviewing for top-tier journals and conferences. He has received the Outstanding Contribution Award at the IJCAI Workshop and the Best Paper Award at the IROS RODGE Workshop.
[Uncaptioned image] Naixin Zhai received the B.S. degree from Hefei University of Technology, Hefei, China, in 2024. He is current working toward the M.S. degree with the Department of Electronic Engineering and Information Science. His research interests include deep learning and synthetic image detection.
[Uncaptioned image] Changtao Miao received his B.S. degree in 2019 from AnHui University. He is currently pursuing the Ph.D. degree in Cyber Science and Technology in University of Science and Technology of China. His research interests include face forgery forensics and face manipulation.
[Uncaptioned image] Zitong Yu (Senior Member, IEEE) received the Ph.D. degree in computer science and engineering from the University of Oulu, Finland, in 2022. Currently, he is an Assistant Professor with Great Bay University, China. He was a Post-Doctoral Researcher with the ROSE Laboratory, Nanyang Technological University. He was a Visiting Scholar with TVG, University of Oxford, from July 2021 to November 2021. His research interests include human-centric computer vision and biometric security. He was a recipient of IAPR Best Student Paper Award, IEEE Finland Section Best Student Conference Paper Award 2020.
[Uncaptioned image] Xingxing Wei received his Ph.D degree in computer science from Tianjin University, and B.S. degree in Automation from Beihang University (BUAA), China. He is now an Associate Professor at Beihang University (BUAA). His research interests include computer vision, adversarial machine learning and its applications to multimedia content analysis. He is the author of referred journals and conferences in IEEE TPAMI, TMM, TCYB, TGRS, IJCV, PR, CVIU, CVPR, ICCV, ECCV, ACMMM, AAAI, IJCAI etc.
[Uncaptioned image] Xun Yang (Senior Member, IEEE) received the Ph.D. degree from the Hefei University of Technology, Hefei, China, in 2017. He is currently a Professor with the Department of Electronic Engineering and Information Science, University of Science and Technology of China (USTC). From 2015 to 2017, he visited the University of Technology Sydney (UTS), Australia, as a Joint Ph.D. Student. He was a Research Fellow with the NExT++ Research Center, National University of Singapore (NUS), from 2018 to 2021. His current research interests include information retrieval, cross-media analysis and reasoning, and computer vision. He serves as an Associate Editor for IEEE TRANSACTIONS ON BIG DATA and Multimedia Systems journal.
[Uncaptioned image] Meng Wang (Fellow, IEEE) is a professor at Hefei University of Technology, China. His current research interests include multimedia content analysis, computer vision, and pattern recognition. He received paper prizes or awards from ACM MM 2009 (Best Paper Award), ACM MM 2010 (Best Paper Award), MMM 2010 (Best Paper Award), SIGIR 2015 (Best Paper Honorable Mention), IEEE TMM 2015 and 2016 (Prize Paper Award Honorable Mention) and ACM TOMM 2018 (Nicolas D. Georganas Best Paper Award), etc. He currently serves on the editorial/advisory boards of IEEE TPAMI, IEEE TMM, IEEE TNNLS, etc. He is a fellow of IEEE and IAPR.
Vulnerabilities in AI-generated Image Detection:
The Challenge of Adversarial Attacks
—Supplementary Document—
Figure 10: Visualization of the spectrum saliency map with different models.
VI Additional Experiments
VI-A Experiment Details
We compared a series of distinguished backbone architectures pretrained on the ImageNet dataset. Specifically, we evaluated ResNet-50, DenseNet-121, EfficientNet-B4, MobileNet-V2, ViT-B/16, and Swin-B. Additionally, we examined Spec equipped with ResNet-34 architecture, as well as DCTA equipped with ResNet-50 architecture. In Table XVIII, we enumerate the accuracy of all models on the test set as observed in our experiments. The meticulous evaluation of these foundational models not only revealed the potential vulnerabilities of AIGI detection models but also highlighted the adversarial threats they face. Our analytical results offer significant insights for understanding and improving the robustness of AIGI detection models.
TABLE XVIII: The accuracy of each detector on different datasets.
Model ResNet DenNet EffNet MobNet Spec DCTA ViT Swin
LSUN(ProGAN) 99.99 99.91 99.96 99.93 99.90 99.91 99.74 100.00
GenImage(SD) 99.55 99.92 99.9 99.93 99.12 99.17 99.48 99.97
FFHQ(Style GAN) 99.81 99.89 99.83 99.63 96.51 96.6 99.55 99.99
VI-B Evaluation on ResNeXt
We use ResNext [xie2017aggregated] as the surrogate to evaluate the attack performance. As reported in Table XIX, FPBA still achieves the best average attack success rate, which is consistent with other surrogate models.
TABLE XIX: Attack evaluation using ResNeXt as surrogate model on the ProGAN-LSUN dataset.
Attack ResNeXt ResNet DenNet Spec DCTA ViT Swin Ave.
MI-FGSM 54.2 29.7 21.3 24.8 49.7 14.2 6.6 28.6
PGD 77.1 35.4 30.9 28.0 50.4 20.6 28.0 38.6
S2I 98.6 90.6 86.9 13.4 36.9 15.0 68.0 58.5
FPBA 99.9 89.0 89.2 21.5 49.0 29.6 64.1 63.2
VI-C Study of Attack Strength
As shown in Table XX, we report the comparison of the attack success rates of our FPBA with existing methods under different attack strengths. The results reflect a pattern consistent with when $\epsilon$=8/255, where our method can achieve globally optimal results compared to previous methods, and a larger perturbation budget consistently enhances adversarial transferability across various surrogate models. However, as the attack strength increases, the noise added causes more and more damage to the image, and the visual impact becomes greater, which is revealed in Fig. 11.
VI-D Spectrum Saliency Map
In Fig. 10, we present the spectrum saliency map of all detectors for the original samples and the spectrum saliency map after frequency domain transformation.
TABLE $X$: The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN(ProGAN) with $\epsilon$=4/255, 8/255 or 12/255. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
IFGSM 52.1 44.6 14.0 36.8 15.0 41.2 12.9 24.6 27.0
MIFGSM 52.1 50.0 23.5 45.4 15.0 45.1 22.9 38.3 34.3
ResNet-50 PGD 53.5 46.3 15.5 38.7 20.3 44.4 12.4 26.0 29.1
S2I 97.1 68.1 21.1 33.0 7.0 18.3 2.0 17.9 23.0
$\epsilon$=4/255 FPBA(Ours) 98.9 93.7 55.0 79.8 15.8 40.8 18.6 75.1 54.1
IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
ResNet-50 PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
$\epsilon$=8/255 FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
IFGSM 52.1 51.9 28.7 50.0 18.3 49.1 40.6 50.2 41.2
MIFGSM 52.1 51.8 28.7 47.4 37.2 50.0 34.7 38.9 41.2
ResNet-50 PGD 96.3 96.1 67.0 91.9 30.6 48.8 58.6 87.5 68.6
S2I 97.3 88.4 78.5 89.3 28.5 47.6 25.4 85.3 63.3
$\epsilon$=12/255 FPBA(Ours) 98.9 99.3 82.9 98.5 20.7 49.2 68.7 98.0 73.9
IFGSM 4.0 5.6 12.5 75.4 12.3 24.6 4.8 3.4 9.6
MIFGSM 9.0 11.5 16.5 75.4 13.7 26.4 6.4 10.9 13.5
MobileNet-V2 PGD 3.4 5.0 11.8 87.7 15.1 25.7 4.5 3.2 9.8
S2I 6.0 7.0 19.2 93.6 5.9 8.0 1.2 5.6 7.6
$\epsilon$=4/255 FPBA(Ours) 7.9 11.8 26.0 99.5 14.4 25.8 4.5 9.5 14.3
IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
MobileNet-V2 PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
$\epsilon$=8/255 FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
IFGSM 18.8 21.4 27.7 75.4 18.0 40.2 11.4 23.0 22.9
MIFGSM 25.2 25.2 33.7 75.4 22.5 48.7 21.6 25.2 28.9
MobileNet-V2 PGD 34.8 39.8 27.6 99.8 35.4 48.5 14.7 45.1 35.1
S2I 28.1 33.5 44.0 98.6 23.3 34.5 7.9 35.5 29.5
$\epsilon$=12/255 FPBA(Ours) 42.3 45.7 56.3 99.9 27.0 42.0 18.7 49.0 40.1
Figure 11: The visual results of adversarial examples with different perturbation budgets.

### v6

@@ IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE V: Transfer-based attack against SOTA detectors on Synthetic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [54].
@@
Attacks against Deepfake detectors are similar to our attacks. However, Deepfakes mainly manipulate or synthesize faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [29], removing facial forgery traces [22, 56] or simply updating perturbations in the frequency domain to improve attack imperceptibility [25] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [22], TraceEvader [56] and FaceAttacker [25], which are the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table III, FPBA consistently achieves the highest attack success rates across all detectors and datasets, outperforming prior Deepfake-based attacks by margins ranging from 11.5% to 87.5%. FakePolisher performs better than TraceEvader and FaceAttack (albeit still less effective than FPBA), but the former largely reduces the image quality, which is very visible and raises suspicion, as shown in Table XIV and Fig. 8. This is likely because FakePolisher is primarily trained on the face manifold, limiting its ability to reconstruct fine-grained, identity-preserving details in AIGI images.
@@ IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [10] and SVRE [58], i.e., utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table IV. First, compared with setting (1), the heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial examples only generated by CNNSpot with FPBA get a success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
@@ IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [33] extracts the global texture representation; UnivFD [43] captures semantic feature using pre-trained CLIP [44]; LNP [31] extracts the noise pattern of images; LGrad [51] extract gradient information through a pre-trained model; DNF [62] leverages diffusion noise features, extracting them through an inverse diffusion process [5]; FreqNet [50] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [9] introduces a frequency-domain mask strategy for data augmentation during training. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table V. Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability between each other. Nevertheless, FPBA still achieves the best adversarial transferability.
@@
Very recently, reconstruction error-based methods, such as DIRE [55] and AEROBLADE [46] have emerged as competitive approaches for detecting diffusion-generated images. Therefore, we adopt them as victim models in our evaluation. As both methods rely on diffusion models, their performance degrades considerably on GAN-generated data. Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX. We follow [46] to report detection performance using average precision (AP). FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
@@ IV-C2 Attack Cross-Generator Image Detection
One important real-world detection problem is cross-generator image detection, i.e., identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 [47] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [41], SD V1.4 [47], SD V1.5 [47], ADM [5], Wukong [57] and BigGAN [2]. Because Swin-ViT achieves the SOTA results on different subsets [66], we use it as the surrogate model.
@@ IV-D Adversarial Transferability Analysis
Next, considering that the AIGI detection task often suffers from imbalanced recognition accuracy between real and fake samples (as shown in Table VI), we further investigate the attack success rates of FPBA on real images and fake images separately, as shown in Table X. Low-level feature detectors, including FreqMask, FreqNet, GramNet, LGrad, mainly rely on frequency, texture, or spectral cues. We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples. We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [59]. As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack. In contrast, high-level detectors like UnivFD, which rely on global semantic features, exhibit overall lower transferability. This suggests that models leveraging generalizable semantic representations are inherently more robust. Interestingly, we observe that in UnivFD, fake images are more resistant to adversarial attacks than real ones. We will leave the phenomenon to explore in future work.
@@ IV-E Evaluation against Defense Models
Next, we investigate the attack performance against defense models. As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors. In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [36] and TRADES [60]. However, we found that the training process of AT fails to converge. Both AT methods cause a substantial drop in clean accuracy, often to around 50%, which is equivalent to random guessing. Robust accuracy against PGD attacks also remains low, in some cases collapsing to nearly 0%. This phenomenon can be attributed to the intrinsic data imbalance in AIGI detection: models tend to overfit to low-level synthetic artifacts early in training. When AT is applied, these brittle features collapse quickly, and adversarial examples easily cross the decision boundary for both real and fake classes. As a result, the model fails to learn discriminative and robust representations, producing near-random predictions on clean data and negligible robustness gains.
@@
Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies. Considering that AIGI detection typically [54] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [54, 66] to utilize Gaussian blurring and JPEG compression as a defense. In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is also crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 7. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under real-world scenarios.
Next, we follow the suggestions from [54, 66] to simultaneously utilize Gaussian blurring and JPEG compression as a defense. As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@ IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have a higher success rate [10]. It indicates the possibility of gradient masking [1]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 9 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
@@ IV-G2 Evaluation on Deepfake Datasets
As we mentioned before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [48], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [27]. We report the results in Table XIII. Although FPBA is not specifically designed for Deepfake detection, it still achieves the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.